OpenStudy (ibbutibbu.):

Please help! Will reward medals! After a discount of 20%, the price of fitness equipment, including tax, was $73.53. The following equation can be used to find x, the price of the equipment without the discount: x - 0.2x + 0.07(x - 0.2x) = 73.53 What is the price of the fitness equipment without the discount?

OpenStudy (anonymous):

asdsadasdasdsadsadsadadadssadsadsdsadsadsaddsaddasdasd why do you hate my mom

OpenStudy (ibbutibbu.):

?...

OpenStudy (anonymous):

let before discount the original price be x so 20% of x =73.53 20100×x=73.53→x=73.53×10020=73.53×5=367.65 Thus, before discount the original price = 367.65 dollars Now you substitute x= 367.65 in the given eq and you will find your required Ans.

OpenStudy (anonymous):

you had texted me it in my box why??

OpenStudy (ibbutibbu.):

i have no clue what u are talking about, @ShreeShree1501

OpenStudy (anonymous):

@Ibbutibbu. hope that helps

OpenStudy (ibbutibbu.):

Thank you, @Lovee_Stasiaa

OpenStudy (anonymous):

plz stay out of my chat box and saying bad things about m y mom

OpenStudy (ibbutibbu.):

please go away, @ShreeShree1501

OpenStudy (anonymous):

bye hahahhahahahaah it was a joke cya

OpenStudy (ibbutibbu.):

ok lol